# PartSDF: Part-Based Implicit Neural Representation for Composite 3D Shape Parametrization and Optimization

### v2

@@ E.2 Shape Generation
E.3 Shape Manipulation
Finally, we perform more manipulation in Figure˜14 with PartSDF on all dataset. By changing part latents, the affected parts have modified appearances but adapt to the original pose and to each others. For example, the car bodies have the same general scales but their wheel wells fit correctly the original wheels, the mixers helices fit the original tube, and the chairs keep the same overall structure while each parts are locally different. When editing their poses, the parts are moved and deformed, but also preserve their original features, e.g., the cars keep their original shapes but adapt to their new wheels, the mixers and their helices are rescaled but conserve their inside/outside relationship, and the size and height of chairs can be changed while maintaining the chairs’ features. Note that this requires the part poses to be correct and coherent. Indeed, if they are manually set to be out of contact and disorganized in 3D space, the resulting shape will be meaningless.

### v3

@@ G.2 Shape Generation
G.3 Shape Manipulation
Finally, we perform more manipulation in Figure˜21 with PartSDF on all dataset. By changing part latents, the affected parts have modified appearances but adapt to the original pose and to each others. For example, the car bodies have the same general scales but their wheel wells fit correctly the original wheels, the mixers helices fit the original tube, and the chairs keep the same overall structure while each parts are locally different. When editing their poses, the parts are moved and deformed, but also preserve their original features, e.g., the cars keep their original shapes but adapt to their new wheels, the mixers and their helices are rescaled but conserve their inside/outside relationship, and the size and height of chairs can be changed while maintaining the chairs’ features. Note that this requires the part poses to be correct and coherent. Indeed, if they are manually set to be out of contact and disorganized in 3D space, the resulting shape will be meaningless.
@@ G.3 Shape Manipulation
Appendix H Single-View Reconstruction
While our main application focuses on shape optimization, we also experiment with single-view 3D reconstruction to demonstrate the feasibility of integrating PartSDF into an image-based pipeline. For this proof-of-concept, we trained a small image encoder to predict the part latents and poses corresponding to our pretrained part decoder. The encoder is a ResNet-18 model trained on synthetic renderings from ShapeNet (Chang et al., 2015) using the dataset of 3D-R2N2 (Choy et al., 2016a).
In Figure˜22, we show reconstruction examples on held-out test images. Despite the simplicity of this setup and limited training data, the results demonstrate that PartSDF can, in principle, recover coherent part-aware 3D structures from single RGB images. Extending this to real-image datasets is a promising direction for future work.
Figure 22: Single-view reconstruction. Given a single RGB image (top), a lightweight ResNet-18 encoder predicts part latents and poses for our pretrained part decoder, which reconstructs the composite 3D shape (bottom). The model recovers coherent part layouts and overall geometry.